OpenStudy (nerdsarecool):

Change the fractions to percents using a strategy of your choice A. 13/20 B. 7/10 C. 12/25 D. 9/30

OpenStudy (nerdsarecool):

Some body please help me

OpenStudy (ben00):

this is very easy. just multiply to get 100.

OpenStudy (ben00):

For example 7/10 actually isn't neccary.

OpenStudy (nerdsarecool):

65 percent correct]

OpenStudy (nerdsarecool):

I just like checking my answers

OpenStudy (ben00):

You can make it easy. By the way it look easy by looking at it. Like 7/10 is 70%

OpenStudy (ben00):

By making it easy i meant by maybe multiplying 13/20 by 5.

OpenStudy (nerdsarecool):

Ok

OpenStudy (ben00):

You would get 65/100 and that's an easy answer.

OpenStudy (ben00):

Or 12/25 times 4 = 48/100. Just like that.

OpenStudy (ben00):

Medal me if you understand.

OpenStudy (danjs):

Given fraction is what over 100? given - 9/30 is - = what over 100 - x/100 \[\frac{ 9 }{ 30 }=\frac{ x }{ 100 }\]

OpenStudy (ben00):

9/30 is tricky but you can always make it easy. By changing it to 3/10.

OpenStudy (danjs):

percent means, out of 100 9/30 is the same as X over 100

OpenStudy (ben00):

You can multiply it. 3/10 x 10 = 30/100 You get 30%

OpenStudy (danjs):

\[\frac{ 9 }{ 30 }=\frac{ x }{ 100 }\] 9*100 = 30*x 900 = 30x x = 30 There you get 30/100 or 30% same for any others, fraction = x/100

OpenStudy (ben00):

I think you can do that to.

OpenStudy (ben00):

Can I get a medal?

OpenStudy (ajspeller):

Just divide the fraction and multiply by 100.